HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

AB INITIO LINE LISTS FOR FOUR ISOTOPOLOGUES OF TRIPLET (X̃^{3}B_{1}) METHYLENE

OLEG EGOROV, Laboratory of Theoretical Spectroscopy, V.E. Zuev Institute of Atmospheric Optics SB RAS 1, Akademician Zuev Sq., Tomsk, 634055 Russia, Faculty of Physics Tomsk State University 36, Lenin Ave., Tomsk, 634050 Russia;

\underline{MICHAËL REY} , \underline{DOMINIKA VIGLASKA} , Groupe de Spectrométrie Moléculaire et Atmosphérique UMR CNRS 7331, UFR Sciences BP 1039, 51687

Reims Cedex 2, France; ANDREI V. NIKITIN, Laboratory of Theoretical Spectroscopy, V.E. Zuev Institute of Atmospheric Optics SB RAS 1, Akademician Zuev Sq., Tomsk, 634055 Russia, Faculty of Physics Tomsk State University 36,

Lenin Ave., Tomsk, 634050 Russia

Complete line lists for four isotopologues of methylene (CH_{2}, CHD, CD_{2}, and 13CH_{2}) in its triplet electronic state (X̃^{3}B_{1}) have been variationally computed using accurate ab initio potential energy and dipole moment surfaces (PES and DMS). In the triplet state, methylene exhibits a large-amplitude bending vibration and can reach a quasilinear configuration due to its low barrier (∼2000 cm^{−1}), making necessary to use a suitable nuclear-motion Hamiltonian^{1}. For the first time, the predicted ab initio height of the barrier to linearity (1924.6 cm^{−1}) agrees well with that obtained from an empirical PES published almost 40 years ago (1931±30 cm^{−1})^{2}. The available experimental band origins for the three isotopologues (CH_{2}, CD_{2}, and ^{13}CH_{2}) are reproduced with errors less than 0.1 cm^{−1} while the total root mean square deviation with the empirical transitions including those for CHD is of 0.112 cm^{−1}.

Due to the very pronounced centrifugal distortion, the rovibrational energy levels strongly overlap, even at rather small rotational quantum numbers, which renders meaningless the use of standard polyad models. In that case, the effect of the rovibrational interactions cannot be ignored, even for the ground vibrational state (000) with K_{a}≥ 5 and global variational calculations where all couplings are inherently taken into account are preferred. Comparison of our ab initio line positions and intensities with previous experimental works will be given. This work has been submitted for publication^{3}.

This work was supported by the Russian Scientific Foundation (RSF, No. 22-42-09022) and Agence Nationale de la Recherche (ANR, French National Research Agency, Grant No. 21-CE30-0053-01).

1M. Rey, D. Viglaska, O. Egorov, A. V. Nikitin, (2023, submitted).

2doi:10.1063/1.450944, P. R. Bunker, P. Jensen, W. P. Kraemer, R. Beardsworth, J. Chem. Phys., 85, 3724, (1986).

3O. Egorov, M. Rey, D. Viglaska, A. V. Nikitin, (2023, submitted).

p-number: p090  Submitted on Mon, 12 Jun 2023 19:56:35 +0200